# Efficient Egocentric Human Activity Recognition

### v1

@@ Abstract
Egocentric video-based models capture rich semantic information and have demonstrated strong performance in human activity recognition (HAR). However, their high power consumption, privacy concerns, and dependence on lighting conditions limit their feasibility for continuous on-device recognition. In contrast, inertial measurement unit (IMU) sensors offer an energy-efficient and privacy-preserving alternative, yet they suffer from limited large-scale annotated datasets, leading to weaker generalization in downstream tasks. To bridge this gap, we propose COMODO, a cross-modal self-supervised distillation framework that transfers rich semantic knowledge from the video modality to the IMU modality without requiring labeled annotations. COMODO leverages a pretrained and frozen video encoder to construct a dynamic instance queue, aligning the feature distributions of video and IMU embeddings. By distilling knowledge from video representations, our approach enables the IMU encoder to inherit rich semantic information from video while preserving its efficiency for real-world applications. Experiments on multiple egocentric HAR datasets demonstrate that COMODO consistently improves downstream classification performance, achieving results comparable to or exceeding fully supervised fine-tuned models. Moreover, COMODO exhibits strong cross-dataset generalization. Benefiting from its simplicity, our method is also generally applicable to various video and time-series pre-trained models, offering the potential to leverage more powerful teacher and student foundation models in future research. The code is available at https://github.com/Breezelled/COMODO .
1 Introduction
Egocentric human activity recognition (HAR) has gained increasing attention due to its applications in healthcare, fitness tracking, and human-computer interaction. With the increasing availability of smart glasses, headsets, and next-generation wearable devices, these devices often integrate multiple sensors, including cameras and inertial measurement units (IMUs), enabling multimodal sensing for HAR.
@@ 1 Introduction
Figure 1: Motivation: Egocentric videos provide rich semantic information but are impractical for continuous on-device recognition, while IMU sensors are lightweight and energy-efficient yet lack large-scale training data. To bridge this gap, we propose cross-modal, self-supervised distillation to enhance IMU representations by leveraging video knowledge.
@@
Recent advances in large-scale video pretraining [40, 2, 27] have significantly enhanced video-based HAR. The availability of large-scale datasets [5, 3, 17, 18, 19] has fueled the success of deep learning models in this domain. Although the video modality achieves high accuracy in human activity recognition, it is often impractical for lightweight devices due to high power consumption, privacy concerns, varying lighting conditions, frequent interruptions, etc. Unlike cameras, IMUs operate independently of lighting conditions and visual occlusions while consuming significantly less power, making them ideal for always-on, on-device activity recognition in resource-constrained scenarios. This efficiency and robustness make IMUs a compelling alternative for HAR on lightweight devices. Despite these advantages, its available training data is significantly more limited. Unlike fields such as computer vision and natural language processing, which have seen significant advancements driven by large-scale datasets, human activity recognition using wearable sensors still relies on relatively small datasets, many of which are collected in controlled environments with scripted activities [47, 1, 35, 6]. A key challenge is that annotating IMU data is much more labor-intensive and costly than labeling images or text, as it often requires external reference signals and precise temporal alignment [21]. This lack of diverse real-world data makes it challenging for data-hungry deep learning models to learn robust representations, limiting their generalization ability in practical applications.
Given the limitations of IMU data, an ideal solution is to leverage the semantic richness of large-scale video training data and pretrained video models while maintaining the efficiency and privacy advantages of IMU sensors. However, there are several challenges to bridge the latent representation from the video domain to IMU domain. Firstly, we need an effective technique to effectively distill a superior video-domain knowledge to IMU domain. Secondly, we need a structured approach to align heterogeneous video and IMU modalities, given their distinct temporal resolutions and feature spaces. Thirdly, given the diversity of pretrained models and downstream tasks, the method needs to be general enough, adaptable to different video and IMU encoder pairs. Finally, the proposed method needs to be both data-efficient (self-supervised) and effectively leverage video representations to enhance IMU-based recognition.
We present CrOss-MOdal video-to-imu DistillatiOn (COMODO) for efficient egocentric HAR, a novel self-supervised learning framework designed to transfer knowledge from the video to IMU modality. Inspired by recent works in self-supervised learning [23], knowledge distillation [10, 29], and multimodal learning [31], our method maintains an instance queue to provide a consistent and large-scale distribution from teacher modality, facilitating the knowledge distillation between IMU and video features.
The main goal of cross-modal knowledge distillation is to leverage knowledge from data-rich modalities to enhance representation learning in data-scarce domains. We demonstrate that COMODO, by performing cross-modal self-supervised distillation, enables pretrained time-series model to inherit rich semantic information from video representations, significantly mitigating the challenges of data scarcity and annotation costs in IMU-based HAR. Our experiments across three benchmark datasets show that COMODO achieves performance comparable to or surpassing fully supervised fine-tuned time-series models, while maintaining the efficiency and privacy advantages of IMU sensors. These results suggest that cross-modal self-supervised learning can bridge the gap between high-resource and low-resource sensing modalities, providing a scalable and effective alternative for real-world wearable AI applications. Our contributions are summarized as follows: (1) Self-supervised Cross-modal Knowledge Transfer: We propose COMODO, a cross-modal self-supervised distillation framework that leverages pretrained video and time-series models enabling label-free knowledge transfer from a stronger modality (video) with richer training data to a weaker modality (IMU) with limited data. (2) A Self-supervised and Effective Cross-modal Queuing Mechanism: We introduce a cross-modal FIFO queue that maintains video embeddings as a stable and diverse reference distribution for IMU feature distillation, extending the instance queue distribution learning approach from single-modality to cross-modality. (3) Teacher-Student Model Agnostic: COMODO supports diverse video and time-series pretrained models, enabling flexible teacher-student configurations and future integration with stronger foundation models. (4) Cross-dataset Generalization: We demonstrate that COMODO maintains superior performance even when evaluated on unseen datasets, and more superior than fully supervised models, highlighting its robustness and generalizability for egocentric HAR tasks.
2 Related Works
Figure 2: Overview of our cross-modal self-supervised distillation framework. The video encoder is pretrained and kept frozen, while the IMU encoder, initialized from a pretrained time-series model, is trained by minimizing the cross-entropy loss between the similarity distributions of video and IMU embeddings, which are computed based on a continuously updated instance queue.
Human Activity Recognition (HAR) has been explored across both vision and sensor-based modalities, each offering unique advantages. Self-supervised learning (SSL) has emerged as a key approach in HAR for leveraging unlabeled sensor data, with contrastive learning effectively capturing generalizable features. [8, 9] introduced contrastive cross-modal learning frameworks for multi-sensor time-series, with [9] improving upon [8] by replacing explicit negative pairs with latent masking. [31] aligned IMU data with vision-language models [34] using contrastive loss, while [39] projected IMU embeddings into a video-based semantic space to enable zero-shot activity recognition. In contrast, our approach performs cross-modal self-supervised distribution distillation from video to IMU.
@@ 2 Related Works
Knowledge distillation (KD) traditionally transfers knowledge from large teachers to smaller students. [38] proposes contrastive representation distillation to align teacher-student embeddings. Similarly, [37, 32] distill rich video representations into IMU models, with [37] needing video at inference and [32] requiring target data fine-tuning. [33] employs multimodal distillation, transferring knowledge from RGB, optical flow, and audio features to a single RGB model, demonstrating that multimodal supervision can improve egocentric HAR. [14] explore mutual distillation, where modality-specific models collaboratively enhance their representations in a supervised setting. For SSL, [10, 29] utilize self-supervised teachers to improve smaller models using the similarity distribution KD, facilitating the generalization of representation in visual and textual domains. [4] frames SSL as self-distillation with a momentum teacher. [43] shows that multimodal students can outperform unimodal teachers by learning richer cross-modal representations. Successful KD depends on the alignment of modality-general, task-relevant features [44].
@@
Unlike [37, 32], which require either video input at inference or fine-tuning on the target dataset, our approach performs self-supervised cross-modal distillation, aligning sensor and video representations without labeled data. While prior self-supervised KD methods [10, 29, 4] focus on single-modal representation learning, our method explicitly transfers knowledge across heterogeneous modalities.
@@
Knowledge transfer between modalities has also been explored using synthetic IMU generation. [25, 26] synthesized IMU signals from video and language models, demonstrating the potential for large-scale HAR by alleviating the need for extensive real-world IMU data collection. More recently, [28] aligned time-series sensor data with large language models, enabling HAR via natural language. These works focus on data synthesis and language-based IMU transfer. Multimodal HAR methods often integrate vision, IMU, and other sensors to improve recognition. [48] applied masked autoencoders for joint representation learning of IMU and video, while [13] introduced temporal attention to optimize multimodal feature integration. [30] proposed a self-supervised framework designed for multi-device HAR that learns robust sensor embeddings via spatial-temporal masking. [24] introduced hierarchical self-supervised pretraining to improve cross-dataset generalization, and [15] handled missing modality conditions through a combination of modality dropout training and cross-modal contrastive learning. To address domain adaptation in egocentric action recognition, [22] applied multimodal few-shot learning. These works reflect the trend of combining multimodal inputs (video, sensor, and language) to overcome data scarcity and domain shifts.
@@
3 Methodology
@@
As discussed in Section 1, while video-based models have demonstrated strong performance in HAR, they come with several practical limitations. On the other hand, IMU sensors are lightweight, power-efficient, and privacy-preserving, making them an attractive choice for real-world applications. However, IMU models lack access to large-scale labeled datasets, making it challenging to learn robust representations. To bridge this gap, we present our CrOss-MOdal self-supervised DistillatiOn (COMODO) that transfers knowledge from the video modality (which benefits from large-scale pretraining) to the IMU modality without requiring labeled data, as illustrated in Figure 2.
Contrastive SSL relies on a sufficiently large number of negative samples to learn discriminative representations effectively. However, maintaining an extremely large batch size is computationally prohibitive. To balance this trade-off, we adopt a FIFO queue approach [23]. This enables the model to maintain a large and diverse set of negative samples to learn from a much broader distribution of video embeddings while keeping the computational cost low.
@@
We define our dataset as $D=\{(v_{i},x_{i})\}_{i=1}^{|D|}$, where each sample consists of a video window $v_{i}$ and its corresponding IMU window $x_{i}$ The dataset contains $|D|$ paired samples collected from egocentric wearable devices. Our objective is to learn an IMU encoder $f_{x}$ that maps an IMU window $x_{i}$ to a feature representation $z_{x}$ , such that the learned representations capture semantic information from the video modality without requiring labeled data.
3.1 Cross-modal Self-supervised Distillation

### v2

@@ Abstract.
The goal of creating intelligent, human-centered wearable systems for continuous activity understanding faces a fundamental trade-off: Egocentric video-based models capture rich semantic information and have demonstrated strong performance in human activity recognition (HAR), but their high power consumption, privacy concerns, and dependence on lighting limit their feasibility for continuous on-device recognition. In contrast, inertial measurement unit (IMU) sensors offer an energy-efficient, privacy-preserving alternative, yet lack large-scale annotated datasets, leading to weaker generalization. To bridge this gap, we propose COMODO, a cross-modal self-supervised distillation framework that transfers semantic knowledge from video to IMU without requiring labels. COMODO leverages a pretrained and frozen video encoder to construct a dynamic instance queue to align the feature distributions of video and IMU embeddings. This enables the IMU encoder to inherit rich semantic structure from video while maintaining its efficiency for real-world applications. Experiments on multiple egocentric HAR datasets show that COMODO consistently improves downstream performance, matching or surpassing fully supervised models, and demonstrating strong cross-dataset generalization. Benefiting from its simplicity and flexibility, COMODO is compatible with diverse pretrained video and time-series models, offering the potential to leverage more powerful teacher and student foundation models in future ubiquitous computing research. The code is available at this repository: https://github.com/cruiseresearchgroup/COMODO.
Human Activity Recognition, Ubiquitous Computing, Cross-modal Knowledge Distillation, Cross-dataset Generalization, Multi-modal Learning, Self-supervised Learning, Wearable AI, Human-centered Computing
copyright: ccjournal: IMWUTjournalyear: 2026journalvolume: 10journalnumber: 2article: 33publicationmonth: 6doi: 10.1145/3810218ccs: Human-centered computing Ubiquitous computingccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Information systems Spatial-temporal systemsccs: Computing methodologies Computer vision representationsccs: Computing methodologies Unsupervised learning
1. Introduction
@@
Figure 1. Motivation: Egocentric videos provide rich semantics but are impractical for continuous on-device recognition, while IMU sensors are lightweight and energy-efficient yet lack large-scale training data. To bridge this gap, we propose cross-modal, self-supervised distillation to enhance IMU representations by leveraging video knowledge.
The vision of human-centered computing hinges on creating technology that seamlessly understands and augments human experiences. A cornerstone of this vision is robust Egocentric Human Activity Recognition (HAR), which has profound applications in healthcare, fitness tracking, and human-computer interaction. The increasing availability of smart glasses, headsets, and next-generation wearables, which often integrate multiple sensors like cameras and inertial measurement units (IMUs), provides an opportunity to realize this vision through rich, multimodal sensing.
@@
Recent advances in large-scale video pretraining (Tong et al., 2022; Bertasius et al., 2021; Li et al., 2023) have significantly enhanced video-based HAR. The availability of large-scale datasets (Carreira and Zisserman, 2017; Caba Heilbron et al., 2015; Goyal et al., 2017; Grauman et al., 2022, 2024) has fueled the success of deep learning models in this domain. Although the video modality achieves high accuracy in human activity recognition, it is often impractical for lightweight devices due to high power consumption, privacy concerns  (Bhardwaj et al., 2024; Mujirishvili et al., 2024; Liao et al., 2025; Opaschi and Vatavu, 2020), varying lighting conditions, frequent interruptions, etc. Unlike cameras, IMUs operate independently of lighting conditions and visual occlusions while consuming significantly less power, making them ideal for always-on, on-device activity recognition in resource-constrained scenarios. This efficiency and robustness make IMUs a compelling alternative for HAR on lightweight devices. Despite these advantages, its available training data is significantly more limited. Unlike fields such as computer vision and natural language processing, which have seen significant advancements driven by large-scale datasets, human activity recognition using wearable sensors still relies on relatively small datasets, many of which are collected in controlled environments with scripted activities (Zhang and Sawchuk, 2012; Anguita et al., 2013; Reiss and Stricker, 2012; Chavarriaga et al., 2013). A key challenge is that annotating IMU data is much more labor-intensive and costly than labeling images or text, as it often requires external reference signals and precise temporal alignment (Haresamudram et al., 2025). This lack of diverse real-world data makes it challenging for data-hungry deep learning models to learn robust representations, limiting their generalization ability in practical applications.
Given the limitations of IMU data (see Figure 1), a promising direction is to leverage the semantic richness of large-scale video training data and pretrained video models while maintaining the efficiency and privacy advantages of IMU sensors. However, there are several challenges to bridge the latent representation from the video domain to IMU domain: (i) Knowledge transfer. Designing objectives and optimization to reliably transfer semantic superior video domain knowledge to IMU domain, which has a significant modality and performance gap (Gong et al., 2023). (ii) Cross-modal alignment. Aligning heterogeneous video and IMU signals with different temporal resolutions and feature spaces (Dhekane and Ploetz, 2025). (iii) Generalization. Ensuring robustness and generalization across various encoder architectures, activity classes, and sensing devices. (iv) Data efficiency. Achieving strong IMU representations with minimal labels via self-supervised signals and scalable training.
Methods Inf. Modality Paradigm Egocentric Self-supervised Structure
VideoMAE (Tong et al., 2022) Video Masked Rec. -
TimeSformer (Bertasius et al., 2021) Video Cross-Entropy -
DeepConvLSTM (Ordóñez and Roggen, 2016) IMU Cross-Entropy -
Attend (Abedin et al., 2021) IMU Cross-Entropy -
CrossHAR (Hong et al., 2024) IMU Masked Rec. + Contrastive -
IMUGPT 2.0 (Leng et al., 2024) IMU Synthetic Generation -
MuJo (Fritsch et al., 2025) IMU Synthetic + Contrastive
IMU2CLIP (Moon et al., 2023) IMU Contrastive
COMODO (Ours) IMU Distribution Distillation
Table 1. Comparison of COMODO with representative HAR frameworks. Inf. Modality denotes the modality used for inference. Paradigm refers to the core learning objective, where “Rec.” stands for Reconstruction and “Contrastive” denotes standard instance-level contrastive learning (i.e., InfoNCE). Egocentric indicates whether the method is compatible with egocentric data. Structure indicates whether the method preserves the semantic structure from the teacher modality. “-” denotes not applicable.
To tackle the challenges above, we present CrOss-MOdal video-to-imu DistillatiOn (COMODO) for efficient egocentric HAR, a novel self-supervised learning framework. Inspired by recent advances in self-supervised learning (He et al., 2020), knowledge distillation (Fang et al., 2021; Limkonchotiwat et al., 2022), and multimodal representation learning (Moon et al., 2023; Fritsch et al., 2025), COMODO is designed to effectively align modalities that differ significantly in temporal resolution, signal characteristics, and feature spaces. As summarized in Table 1, distinct from prior approaches that rely on heavy video inference (Tong et al., 2022; Bertasius et al., 2021) or instance-wise contrastive alignment (Moon et al., 2023; Fritsch et al., 2025) which neglects the teacher’s semantic structure, COMODO distills structured knowledge by aligning the similarity distribution between video and IMU embeddings. This is facilitated by a dynamic instance queue that provides a stable and diverse reference distribution, introducing a beneficial temporal prior that stabilizes the cross-modal optimization process and enhances representation learning. Practically, COMODO decouples training-time paired video supervision from IMU-only deployment, providing a label-free, model-agnostic way to transfer semantic structure from data-rich video to efficient, privacy-preserving IMU sensing.
Cross-modal distillation aims to leverage a data-rich modality to improve representations in a data-scarce domain. We show that COMODO enables pretrained time-series encoders to inherit semantic information from video representations, mitigating data scarcity and annotation costs in IMU-based HAR. Evaluated across three egocentric datasets comprising over 70 activity categories, including fine-grained datasets with more than 30 distinct daily activities, COMODO matches or surpasses fully supervised fine-tuned time-series models while retaining IMU efficiency and privacy, suggesting a scalable path to real-world wearable AI (Figure 2).
Our contributions are summarized as follows:
  • Self-supervised Cross-modal Knowledge Transfer. COMODO enables label-free distillation from a stronger (video) to a weaker (IMU) modality using pretrained video and time-series models. Leveraging rich pre-trained knowledge from the video domain to alleviate the challenge of scarce data for wearable IMU sensors.
  • Effective Cross-modal Queuing Mechanism. We introduce a cross-modal FIFO queue that maintains video embeddings as a stable, diverse reference distribution for IMU distillation, narrowing the modality gap and handling heterogeneity.
  • Model-agnostic Teacher–Student. COMODO supports diverse pretrained video and time-series models, allowing flexible teacher–student configurations and future integration with stronger foundation models.
  • Cross-dataset Generalization. To validate real-world applicability, we show that COMODO maintains superior performance on unseen datasets with different classes and devices, outperforming supervised models and highlighting robustness and generalizability for egocentric HAR.
Figure 2. Examples of real-world applications of COMODO-enabled egocentric HAR system. Note: Videos are used only during training to transfer rich semantic knowledge, while lightweight IMU sensors are used at inference for efficient, privacy-preserving operation.
2. Related Works
Our work is situated at the intersection of human activity recognition, self-supervised learning, and cross-modal knowledge distillation. To contextualize our contributions, this section reviews prior art from these interconnected domains. We position COMODO by first examining (1) self-supervised learning for HAR, starting from intra-modal sensor-based approaches and extending to cross-modal paradigms. We then discuss (2) the broader field of knowledge distillation, covering its evolution from supervised to self-supervised and cross-modal applications. Finally, we contrast our distillation-based approach with (3) alternative multimodal strategies such as data synthesis and multimodal fusion.
2.1. Human Activity Recognition
Human Activity Recognition (HAR) has been explored across both vision and sensor-based modalities, each offering unique advantages. To mitigate the reliance on large-scale labeled sensor datasets, self-supervised learning (SSL) has emerged as a key approach, with contrastive learning effectively capturing generalizable features. Recent works in sensor-based HAR have successfully leveraged unlabeled data through various self-supervised learning strategies. For instance, some approaches utilize a teacher-student framework to distill knowledge from unlabeled data pools (Tang et al., 2021), while others creatively treat synchronously captured data from multiple on-body devices as natural transformations for contrastive learning (Jain et al., 2022). Deldari et al. (2022, 2024) introduced contrastive cross-modal learning frameworks for multi-sensor time-series, with Deldari et al. (2024) improving upon Deldari et al. (2022) by replacing explicit negative pairs with latent masking. Despite their success, these methods primarily focus on extracting supervisory signals from within the sensor modality itself. Miao et al. (2024) proposed a self-supervised framework designed for multi-device HAR that learns robust sensor embeddings via spatial-temporal masking. Hong et al. (2024) introduced hierarchical self-supervised pretraining to improve cross-dataset generalization. Other works have begun to bridge modalities, seeking to leverage richer semantic sources to improve sensor-based HAR, such as Moon et al. (2023) align IMU data with vision-language models (Radford et al., 2021), or projecting IMU embeddings into a video-based semantic space for zero-shot recognition (Tong et al., 2021). Most recently, MuJo (Fritsch et al., 2025) learns a joint feature space using synthetic IMU and extracted poses derived from exocentric videos, which limits its applicability in egocentric scenarios. In contrast to these, our approach performs cross-modal self-supervised distribution distillation from video to IMU.
2.2. Knowledge Distillation
@@
Knowledge distillation traditionally transfers knowledge from large teachers to smaller students. Tian et al. (2020) proposes contrastive representation distillation to align teacher-student embeddings. Similarly, Tan et al. (2023); Ni et al. (2022) distill rich video representations into IMU models. Radevski et al. (2023) employs multimodal distillation, transferring knowledge from RGB, optical flow, and audio features to a single RGB model, demonstrating that multimodal supervision can improve egocentric HAR. Garcia et al. (2021) explore mutual distillation, where modality-specific models collaboratively enhance their representations in a supervised setting. For self-supervised learning, Fang et al. (2021); Limkonchotiwat et al. (2022) utilize self-supervised teachers to improve smaller models using the similarity distribution knowledge distillation, facilitating the generalization of representation in visual and textual domains. Caron et al. (2021) frames self-supervised learning as self-distillation with a momentum teacher. Xue et al. (2021) shows that multimodal students can outperform unimodal teachers by learning richer cross-modal representations. Successful knowledge distillation depends on the alignment of modality-general, task-relevant features (Xue et al., 2023).
@@
Unlike Tan et al. (2023); Ni et al. (2022), which require either video input at inference or fine-tuning on the target dataset, our approach performs self-supervised cross-modal distillation, aligning sensor and video representations without labeled data. While prior self-supervised knowledge distillation methods (Fang et al., 2021; Limkonchotiwat et al., 2022; Caron et al., 2021) focus on single-modal representation learning, our method explicitly transfers knowledge across heterogeneous modalities.
2.3. Multimodal Learning and Synthetic Data Generation for HAR
@@
Knowledge transfer between modalities has also been explored using synthetic IMU generation. Kwon et al. (2020); Leng et al. (2024) synthesized IMU signals from video and language models, demonstrating the potential for large-scale HAR by alleviating the need for extensive real-world IMU data collection. More recently, Li et al. (2024) aligned motion sensor time-series with large language models for HAR, and (Li et al., 2025) further explored agentic, zero-shot HAR from raw motion time-series. These works focus on data synthesis, language-based or agentic sensor understanding for HAR. Multimodal HAR methods often integrate vision, IMU, and other sensors to improve recognition. Zhang et al. (2024) applied masked autoencoders for joint representation learning of IMU and video, while Gao et al. (2023) introduced temporal attention to optimize multimodal feature integration. Gong et al. (2023) handled missing modality conditions through a combination of modality dropout training and cross-modal contrastive learning. To address domain adaptation in egocentric action recognition, Hatano et al. (2024) applied multimodal few-shot learning. These works reflect the trend of combining multimodal inputs (video, sensor, and language) to overcome data scarcity and domain shifts.
@@
3. Methodology
Figure 3. Overview of our cross-modal self-supervised distillation framework. The video encoder is pretrained and kept frozen, while the IMU encoder, initialized from a pretrained time-series model, is trained by minimizing the cross-entropy loss between the similarity distributions of video and IMU embeddings, which are computed based on a continuously updated instance queue.
@@ 3. Methodology
As discussed in Section 1, while video-based models have demonstrated strong performance in HAR, they come with several practical limitations. On the other hand, IMU sensors are lightweight, power-efficient, and privacy-preserving, making them an attractive choice for real-world applications. However, IMU models lack access to large-scale labeled datasets, making it challenging to learn robust representations. To bridge this gap, we present our COMODO that transfers knowledge from the video modality (which benefits from large-scale pretraining) to the IMU modality without requiring labeled data, as illustrated in Figure 3.
Contrastive self-supervised learning relies on a sufficiently large number of negative samples to learn discriminative representations effectively. However, maintaining an extremely large batch size is computationally prohibitive. To balance this trade-off, we adopt a FIFO queue approach (He et al., 2020). This design serves two critical purposes in our cross-modal distillation context. First, it provides a computationally efficient mechanism to maintain a large and diverse set of video embeddings, forming a stable reference distribution for the IMU student to learn from. Second, and more subtly, the FIFO structure imposes an implicit temporal prior on the optimization process. Unlike random sampling from a static bank, gradual updating of the queue ensures that the target distribution evolves smoothly over training iterations. As we demonstrate in our ablation study (Section 4.6.5), this leads to a more stable and coherent optimization flow, which is particularly beneficial for bridging the significant gap between heterogeneous modalities.
@@
We define our dataset as $D=\{(v_{i},x_{i})\}_{i=1}^{|D|}$, where each sample consists of a video window $v_{i}$ and its corresponding IMU window $x_{i}$ . The dataset contains $|D|$ paired samples collected from egocentric wearable devices. Our objective is to learn an IMU encoder $f_{x}$ that maps an IMU window $x_{i}$ to a feature representation $z_{x}$ , such that the learned representations capture semantic information from the video modality without requiring labeled data.
3.1. Cross-modal Self-supervised Distillation